the height of a triangle is 2 meters less than twice the length of the base. if the sail has an area of 30 square meters, find the length of its base and the height.
Answer by Boreal(15235)

Area=1/2(bh)
b=x
height=2x-2, units meters
30=(1/2)bh
60=bh=x(2x-2)
2x^2-2x=60
2x^2-2x-60=0
x^2-x-30=0
(x-6)(x+5)=30
x=6 as only positive root.
2x-2=10
Their product is 60 m^2, half of which is 30 m^2
The base is 6 m; the height is 10 m.